OpenStudy (anonymous):

Which expression is the simplified form of the sqrt(h^11)

OpenStudy (anonymous):

\(h^{\frac{11}{2}}\)

OpenStudy (anonymous):

is that what you mean?

OpenStudy (anonymous):

well no disregard the question lol, \[\sqrt{h ^{11}}\] what is the simplified version of this

OpenStudy (anonymous):

I have a trick, if you like , you can take it. You see \(\sqrt \) is 1/2 , right? that means 2 is in denominator. You take 11\(\div\) 2 = 5 remainder 1. so, 5 is exponent of h in outside of the squaroot, 1 is exponent of h inside of the square root. That is \(h^5\sqrt {h^1}= h^5\sqrt{h}\) It is always right for any root with any exponent

OpenStudy (anonymous):

oh my god ur amazing lol thank u so much lol

OpenStudy (anonymous):